# INSIGHT: Explainable Weakly-Supervised Medical Image Analysis

### v2

@@ 3.4 Training Objective
Binary Cross-Entropy with Label Smoothing. BCE loss with label smoothing improves generalization and prevents overconfidence in noisy or overlapping labels:
$\mathcal{L}_{\text{BCE}}=-\frac{1}{N}\sum_{i=1}^{N}\left[y^{\prime}_{i}\log(\hat{y}_{i})+(1-y^{\prime}_{i})\log(1-\hat{y}_{i})\right],$ (6)
where $y^{\prime}_{i}=(1-\epsilon)y_{i}+\frac{\epsilon}{2}$ is the smoothed label, $\epsilon$ is the smoothing factor, and $\hat{y}_{i}$ is the predicted probability.
Spectral Decoupling. To address gradient starvation and encourage robust decision boundaries, we include spectral decoupling loss [29]:
@@
where $z$ are unnormalized logits, and $\lambda_{\text{SD}}$ controls regularization strength. Spectral decoupling is one of the most effective methods for overcoming dataset bias without additional sub-group labels [36, 37].
@@ 3.5 Heatmap Visualization
Heatmaps are resized and reconstructed to match the original image dimensions. For WSIs, patches are stitched based on coordinates, while for CT volumes, slices are stacked sequentially. Dynamic thresholding with Otsu’s algorithm [27] isolates regions of interest:
$\mathcal{H^{\prime}}=\mathcal{H}\cdot\mathbb{I}(\mathcal{H}>T),$ (9)
where $T=\arg\min_{t}\sigma_{w}^{2}(t)$ minimizes intra-class variance.
@@ 4 Experiments
4.1 Datasets and Preprocessing
We evaluate INSIGHT on three publicly available datasets representing two imaging modalities: CT volumes and WSIs.
4.1.1 Computed Tomography (CT) Dataset
MosMed [24]. The MosMed dataset comprises $1,110$ chest CT studies for COVID-19 detection, split into two subsets: MosMed-A ($1,060$ volumes) and MosMed-B ($50$ volumes). MosMed-A is used for classification tasks, employing five-fold cross-validation with an 80/20 train-test split. MosMed-B, containing voxel-level annotations, is used solely for evaluating segmentation performance by comparing INSIGHT-generated heatmaps $\mathcal{H}$ against ground truth using Dice score.
CT volumes are normalized to the Hounsfield Unit (HU) range $[-1000,400]$, scaled to $[0,1]$, and resized to $518\times 518\times 32$. To ensure compatibility with pre-trained ViTs, single-channel data is replicated across three channels. Lung parenchyma is isolated using Lungmask [14].
4.1.2 Whole Slide Image (WSI) Datasets
CAMELYON16 [18]. This dataset contains $399$ WSIs for detecting metastatic breast cancer in lymph nodes. We split the $269$ training slides into 90/10 for training and validation. The model checkpoint with the highest validation Dice score is evaluated on the official test set of $129$ slides.
BRACS [1]. This dataset contains $547$ WSIs annotated with $4,539$ bounding boxes for breast carcinoma subtype classification. Following BRACS’s official splits, the final dataset comprises $347$, $51$, and $73$ slides for training, validation, and testing, respectively.
Both CAMELYON16 and BRACS use the CLAM toolbox [21] for patch extraction, preserving spatial embeddings for INSIGHT.
Table 1: Performance comparison on the CAMELYON16 and BRACS official test datasets. ALL aggregators use UNI [8] for embeddings. We report AUC and Dice score (mean ± std) for CAMELYON16, and AUC scores for different categories in BRACS, with Macro AUC as the overall score. Reported Dice scores for all comparison methods were fine-tuned through grid search on the validation set.
Aggregator CAMELYON16 [18] BRACS [1]
AUC Dice(%) ADH FEA DCIS Invasive Macro AUC
ABMIL [15] 0.975 55.8$\pm$25.0 0.656 0.744 0.804 0.995 0.800
CLAM-SB [21] 0.966 64.7$\pm$24.1 0.611 0.757 0.833 0.999 0.800
CLAM-MB [21] 0.973 67.7$\pm$22.6 0.701 0.687 0.828 0.998 0.804
TransMIL [35] 0.982 12.4$\pm$22.4 0.644 0.653 0.769 0.989 0.764
INSIGHT 0.990 74.6$\pm$19.1 0.734 0.790 0.837 0.999 0.840
Table 2: Performance comparison for classification and segmentation on the MosMed dataset. Methods marked with * were trained on an external dataset but tested on the same MosMed test set for consistency.
Task Method Setting AUC / Dice(%)
Classification ViT-COVID [38] 0.6/0.2/0.2 split 0.870
PR-3D-CNN [13] Five-fold CV 0.914 ± 0.049
INSIGHT Five-fold CV 0.962 ± 0.012
Segmentation 3D U-Net [2]* Voxel-level 40.5 ± 21.3
3D GAN [34] Volume-level 41.2 ± 14.7
INSIGHT Volume-level 42.7 ± 15.3
4.2 INSIGHT’s Implementation Details
Architecture. INSIGHT uses ViT-L/DINOv2 for CT slices and UNI for WSI tiles. Pre-processed inputs are resized to $224\times 224$, and embeddings from the last layer of the encoder are concatenated to retain spatial information. Dimensionality is reduced from $1024$ to $128$ using a linear transformation. Detection and Context modules consist of three convolutional layers with GELU activations, layer normalization, and a final layer for label-specific outputs.
Training Configuration. INSIGHT is trained using AdamW with a learning rate of $10^{-4}$ and weight decay of $10^{-4}$. Training is capped at $50$ epochs, with early stopping triggered after $8$ epochs of no validation improvement. Hyperparameters, including $\alpha=8$ for SmoothMax pooling, are tuned on validation sets. Label smoothing is applied for multi-label classification on BRACS, while Spectral Decoupling is used for MosMed and CAMELYON16 to improve generalization.
@@ 4.3 Compared Baselines
We compare INSIGHT to several state-of-the-art methods across the MosMed, CAMELYON16, and BRACS datasets, evaluating both classification and segmentation performance.
MosMed. For the MosMed dataset, we benchmark against methods specifically designed for COVID-19 CT analysis:
  • VIT-COVID Classifier [38]: A ViT pre-trained on a large-scale dataset of COVID-19 CT images, designed for binary classification.
  • Progressively Resized 3D-CNN [13]: A CNN that progressively resizes input volumes to extract multi-scale information, enhancing analysis in volumetric CT images.
  • 3D U-Net [2]: A fully supervised segmentation model trained with voxel-level annotations, providing a direct comparison compared to weakly supervised methods.
  • 3D GAN [34]: A weakly supervised generative model that performs volume-level segmentation without requiring dense voxel-level annotations.
WSI Datasets. For the CAMELYON16 and BRACS datasets, we compare INSIGHT against leading MIL-based methods. To ensure a fair comparison, all methods (including ours) use the UNI encoder [8] for feature extraction:
  • ABMIL [15]: An attention-based MIL approach that identifies the most informative regions within a slide, commonly used in weakly supervised WSI analysis.
  • CLAM-SB/MB [21]: A MIL method combining attention mechanisms with clustering to improve localization. CLAM-SB uses a single branch for non-overlapping classes, while CLAM-MB employs multiple branches to target specific classes.
  • TransMIL [35]: A transformer-based MIL model that captures long-range dependencies within WSIs, enhancing contextual analysis for slide-level classification.
Figure 2: Qualitative comparison of heatmaps generated by INSIGHT (ours) and baseline methods. For CAMELYON16, we compare lesion detection heatmaps generated by (a) our INSIGHT model and (b) the baseline methods. Additionally, we present (c) heatmaps for MosMed and (d) heatmaps for BRACS. Ground-truth areas are outlined in green. Note that BRACS does not provide detailed annotations, so ground truth is not displayed in panel (d). Additional visualizations can be found in the Appendix (Sec. A) for further reference.
@@ 4.4 Quantitative Results
INSIGHT consistently outperformed baseline methods in classification and segmentation tasks across all datasets, demonstrating its adaptability to diverse medical imaging modalities. Performance metrics are reported in Table 1 and Table 2.
Key Results:
  • On CAMELYON16, INSIGHT achieved a Dice score of 74.6%, outperforming the top competing model by 6.9%. This significant improvement highlights its ability to enhance boundary precision in segmentation tasks.
  • On BRACS, INSIGHT demonstrated strong multi-label classification performance, achieving AUC improvements of 3.3% for both ADH and FEA subtyping. This showcases its effectiveness in distinguishing subtle tissue variations critical for breast carcinoma subtype classification.
  • On MosMed, INSIGHT achieved a classification AUC of 96.2%, nearly 5% higher than the best baseline, underscoring its robustness in volumetric CT analysis for COVID-19 detection.
Table 3: Ablation study results on the MosMed, CAMELYON16, and BRACS datasets. INSIGHT performance is shown with different configurations of CS (Context Suppression), SM (SmoothMax), and Rg (Regularizer).
Config MosMed [24] CAMELYON16 [18] BRACS [1]
CS SM Rg AUC Dice(%) AUC Dice(%) ADH FEA DCIS Invasive
0.968$\pm$0.016 34.9$\pm$15.2 0.992 25.5$\pm$16.3 0.741 0.509 0.764 0.993
0.947$\pm$0.032 36.9$\pm$15.7 0.982 42.1$\pm$21.4 0.733 0.507 0.799 0.998
0.955$\pm$0.026 42.9$\pm$15.4 0.969 76.7$\pm$15.7 0.728 0.767 0.834 0.999
0.962$\pm$0.012 42.7$\pm$15.3 0.990 74.6$\pm$19.1 0.734 0.790 0.837 0.999
@@
INSIGHT’s superior performance is attributed to three key architectural innovations:
@@
Spatial Embedding. By retaining spatial resolution during feature extraction, INSIGHT produces feature maps of $16\times 16\times 1024$ for CT and $14\times 14\times 1024$ for WSIs, preserving diagnostically meaningful spatial structures. In contrast, baseline methods reduce feature maps to $1\times 1\times 1024$ vectors via pooling, discarding fine-grained spatial information. This capability enables INSIGHT to detect small but critical patterns, particularly in datasets like BRACS, where distinct subtypes may co-exist within a single slide.
@@
Adjacent Context Focus. INSIGHT’s context module captures localized dependencies through larger convolutional filters, avoiding the noise introduced by global self-attention mechanisms. This focus on adjacent spatial relationships is particularly effective for multi-label datasets like BRACS, where lesion subtypes often appear in close proximity.
@@
Broader Foreground Activation Consideration. Unlike conventional MIL-based models that focus only on the top $k$ patches, INSIGHT’s smooth pooling method considers all relevant foreground activations above a threshold. This approach preserves finer boundary details, significantly enhancing segmentation accuracy. For example, on CAMELYON16, this broader activation inclusion contributed to its 6.9% Dice score improvement.
INSIGHT’s performance improvements are further supported by qualitative results (Fig. 2) and ablation studies (Sec. 4.6), demonstrating the critical role of its architectural innovations in achieving state-of-the-art results.
@@ 4.5 Qualitative Analysis
We visualize and compare heatmaps generated by INSIGHT and baseline methods in Fig. 2. Unlike baseline methods, which often rely on post-hoc adjustments, INSIGHT directly integrates interpretability into its architecture, mapping each patch’s probability to the heatmap. This eliminates the need for additional calibration steps, producing well-calibrated, fine-grained heatmaps that effectively highlight diagnostically significant regions. Additional visualizations are provided in Appendix A, which further illustrate the superior interpretability and precision of INSIGHT’s heatmaps across both WSIs and CT datasets.
@@
CT Heatmap Comparisons:
INSIGHT demonstrates strong generalization across modalities, accurately highlighting infected regions in CT volumes. This is particularly evident in MosMed, where INSIGHT’s heatmaps precisely localize areas of COVID-19 infection with minimal false positives or negatives. These findings underscore INSIGHT’s versatility and robustness compared to baseline methods.

### v3

@@ 3.3 Training Objective
Binary Cross-Entropy Loss. Since our model is trained using only volume- or slide-level categorical labels, we optimize the predicted probabilities using the binary cross-entropy loss $\mathcal{L}_{\text{BCE}}$. For multi-label settings, the loss is independently computed for each class.
Spectral Decoupling. To address gradient starvation and encourage robust decision boundaries, we include spectral decoupling loss (Pohjonen et al., 2022):
@@
where $z$ are unnormalized logits, and $\lambda_{\text{SD}}$ controls regularization strength. Spectral decoupling is one of the most effective methods for overcoming dataset bias without additional sub-group labels (Shrestha et al., 2022a, b).
@@ 3.4 Heatmap Visualization
Heatmaps are reconstructed to match the original image dimensions for intuitive interpretation. For WSIs, patch-level heatmaps are stitched together based on their spatial coordinates. For CT volumes, slice-level heatmaps are stacked in order to reconstruct the full volume. The resulting heatmaps are then upsampled using bicubic interpolation to ensure smooth and visually consistent representations. The pseudocode for the heatmap generation and classification process is provided in Appendix D.
4 Datasets and Preprocessing
We evaluate INSIGHT on three publicly available datasets representing two imaging modalities: whole slide images (WSI) and CT volumes.
4.1 Computed Tomography Dataset
MosMed (Morozov et al., 2020). The MosMed dataset comprises $1,110$ chest CT studies for COVID-19 detection, split into two subsets: MosMed-A ($1,060$ volumes) and MosMed-B ($50$ volumes). MosMed-A is used for classification tasks, employing five-fold cross-validation with an 8080/2020 train-test split. MosMed-B, containing voxel-level annotations, is used solely to evaluate segmentation performance by comparing INSIGHT-generated heatmaps $\mathcal{H}$ against ground truth using Dice score. CT volumes are normalized to the Hounsfield Unit (HU) range [1000,400][-1000,400], scaled to $[0,1]$, and resized to $518\times 518\times 32$. To ensure compatibility with pre-trained ViTs, single-channel data is replicated across three channels. Lung parenchyma is isolated using Lungmask (Hofmanninger et al., 2020).
4.2 Whole Slide Image Datasets
CAMELYON16 (Litjens et al., 2018). This dataset contains $399$ WSIs for detecting metastatic breast cancer in lymph nodes. Following UNI (Chen et al., 2024), we use the $20\times$ magnification slides in our experiments, omitting one unusable slide from the original training set. We split the remaining $269$ training slides into $243$ for training and $26$ for validation. The model checkpoint with the highest validation Dice is evaluated on the official test set of $129$ slides.
BRACS (Brancati et al., 2021). This dataset contains $547$ WSIs annotated with $4539$ bounding boxes for breast carcinoma subtype classification: ADH (Atypical Ductal Hyperplasia – early premalignant change), FEA (Flat Epithelial Atypia – precursor lesion), DCIS (Ductal Carcinoma in Situ – non-invasive cancer), and Invasive (Invasive Carcinoma – malignant stromal invasion). As BRACS provides only the most severe label for each slide, we extracted fine-grained labels using QuPath (Bankhead et al., 2017). Following BRACS’s official splits and excluding slides with incomplete labeling or no valid regions, the final dataset comprises $347$, $51$, and $73$ slides for training, validation, and testing. We downsample all WSIs from $40\times$ to $20\times$ magnification.
We use the CLAM toolbox (Lu et al., 2021) for patch extraction for both CAMELYON16 and BRACS.
@@ 5 Experiments
This section presents the experimental setup and results. We provide additional implementation details, including training configurations and model architecture parameters, in Appendix A.
@@ 5.1 Compared Baselines
We compare INSIGHT to state-of-the-art methods across the MosMed, CAMELYON16, and BRACS datasets, evaluating both classification and segmentation performance under a weakly supervised setting.
CT Volumes. For the MosMed dataset, we benchmark against methods specifically designed for COVID-19 CT analysis.
  • Progressively Resized 3D-CNN (Hasan et al., 2021), a CNN that progressively resizes input volumes to extract multi-scale information.
  • 3D U-Net (Bressem et al., 2021), a fully supervised segmentation model trained with voxel-level annotations that provides a direct comparison compared to weakly supervised methods.
  • 3D GAN (Shabani et al., 2022), a weakly supervised generative model that performs volume-level segmentation without requiring dense voxel-level annotations.
WSI Datasets. To ensure a fair comparison, all methods (including ours) use the same foundational model encoder for feature extraction:
  • ABMIL (Ilse et al., 2018): An attention-based MIL approach that identifies the most informative regions within a slide, commonly used in weakly supervised WSI analysis.
  • CLAM-SB/MB (Lu et al., 2021): A MIL method combining attention mechanisms with clustering to improve localization. CLAM-SB uses a single branch for non-overlapping classes, while -MB employs multiple branches to target specific classes.
  • TransMIL (Shao et al., 2021): A transformer-based MIL model that captures long-range dependencies within WSIs, enhancing contextual analysis for slide-level classification.
  • WiKG (Li et al., 2024): A dynamic graph representation learning framework that models knowledge graphs, implementing a knowledge-aware attention mechanism to capture contextual interactions.
@@ 5.2 Quantitative Results
INSIGHT consistently outperformed baseline methods in classification and segmentation tasks across all datasets, demonstrating adaptability to diverse medical imaging modalities. Performance metrics are reported in Table 1 and Table 2. On MosMed, INSIGHT achieved a classification AUC of 96.2%, nearly 5% higher than the best baseline, underscoring its robustness in volumetric CT analysis for COVID-19 detection. On CAMELYON16, INSIGHT achieved a Dice score of 74.6%, outperforming the top competing model by 6.9%, highlighting its ability to enhance boundary precision in segmentation. On BRACS, INSIGHT showed strong multi-label classification performance, with AUC improvements of 3.3% for both ADH and FEA, showcasing its effectiveness in distinguishing subtle tissue variations critical for subtype classification.
@@
INSIGHT’s superior performance stems from three key architectural innovations:
@@
Spatial Embedding. By retaining spatial resolution during feature extraction, INSIGHT produces $14\times 14\times 1024$ feature maps per WSI tile, preserving diagnostically meaningful structure. In contrast, baseline methods reduce feature maps to $1\times 1\times 1024$ via pooling, discarding fine-grained information. This enables INSIGHT to detect small but critical patterns, especially in BRACS, where distinct subtypes may co-exist within a single slide.
Table 1: Performance comparison for classification and segmentation on the MosMed dataset. Methods marked with * were trained on an external dataset but tested on the same MosMed test set for consistency. Baseline details in Sec. 5.1.
Task Method AUC / Dice (%)
Classification PR-3D-CNN (Five-fold CV) 0.914 ± 0.049
INSIGHT (Five-fold CV) 0.962 ± 0.012
Segmentation 3D U-Net (Voxel-level) 40.5 ± 21.3
3D GAN (Volume-level) 41.2 ± 14.7
INSIGHT (Volume-level) 42.7 ± 15.3
Table 2: Performance comparison on the CAMELYON16 and BRACS official test datasets. ALL aggregators use UNI for embeddings. We report AUC and Dice scores (mean ± std) for CAMELYON16, and AUC scores for different categories in BRACS. Reported Dice scores for all comparison methods were fine-tuned through grid search on the validation set. Baseline details in Sec. 5.1. BRACS’s subtype definitions are provided in Sec. 4.2.
Aggregator CAMELYON16 BRACS
AUC Dice ADH FEA DCIS Invasive
[…]
@@
Adjacent Context Focus. INSIGHT’s context module captures local dependencies using larger convolutional filters. This focus on adjacent spatial relationships is especially effective for multi-label datasets like BRACS, where lesion subtypes often appear in close proximity.
@@
Broader Foreground Activation Consideration. Unlike MIL-based models that focus only on top $k$ patches, INSIGHT’s smooth pooling considers all relevant activations above a threshold. This preserves finer boundary details, enhancing segmentation accuracy. For example, on CAMELYON16, this broader inclusion contributed to its 6.9% Dice improvement.
We further conducted experiments using the Virchow 2 foundational model (Zimmermann et al., 2024), a more recent state-of-the-art pretraining approach for histopathology images, to assess INSIGHT’s performance and robustness against other benchmark methods. As shown in Table 3, INSIGHT again largely outperformed all baselines in segmentation task on the CAMELYON16 dataset. INSIGHT’s performance improvements are further supported by qualitative results (Fig. 4) and ablation studies (Sec. 5.5), demonstrating the critical role of its architectural innovations in achieving state-of-the-art results.
Table 3: Performance comparison on the CAMELYON16 official test dataset using Virchow 2 for embeddings. We report AUC and Dice scores (mean ± std) for CAMELYON16. Reported Dice scores for all comparison methods were fine-tuned through grid search on the validation set. Baseline details in Sec. 5.1.
Aggregator AUC Dice (%)
[…]
Table 4: Performance comparison (Dice, mean ± std) on different lesion sizes. Paired one-tailed permutation hypothesis tests (10,000 iterations) compare each baseline to INSIGHT. Bolded indicates $0.01\leq p<0.05$, and ∗∗ indicates $p<0.01$.
Small Moderate Large
Model Aggregator Dice (%) p-value Dice (%) p-value Dice (%) p-value
UNI INSIGHT $\mathbf{42.8\pm 38.6}$ - $\mathbf{74.6\pm 27.3}$ - $\mathbf{82.1\pm 19.8}$ -
CLAM-MB $20.7\pm 32.2$ ∗∗ $62.2\pm 27.8$ ∗∗ $\mathbf{81.4\pm 19.6}$ $0.40$
CLAM-SB $20.8\pm 32.2$ ∗∗ $62.6\pm 27.4$ ∗∗ $\mathbf{81.2\pm 19.2}$ $0.36$
ABMIL $11.0\pm 25.2$ ∗∗ $51.1\pm 29.8$ ∗∗ $72.4\pm 23.0$ ∗∗
TransMIL $30.6\pm 37.3$ ∗∗ $23.4\pm 32.0$ ∗∗ $34.1\pm 32.3$ ∗∗
WiKG $19.5\pm 31.9$ ∗∗ $58.7\pm 29.1$ ∗∗ $\mathbf{80.3\pm 21.0}$ $0.29$
Virchow2 INSIGHT $\mathbf{55.5\pm 37.0}$ - $\mathbf{78.6\pm 22.1}$ - $\mathbf{81.2\pm 19.7}$ -
CLAM-MB $24.5\pm 33.3$ ∗∗ $60.9\pm 24.7$ ∗∗ $\mathbf{81.4\pm 17.7}$ $0.52$
CLAM-SB $27.7\pm 34.3$ ∗∗ $66.6\pm 22.8$ ∗∗ $\mathbf{84.1\pm 16.5}$ $0.81$
ABMIL $10.8\pm 24.4$ ∗∗ $50.9\pm 30.1$ ∗∗ $74.6\pm 23.6$ $0.04$
TransMIL $42.4\pm 37.0$ ∗∗ $35.4\pm 32.8$ ∗∗ $50.9\pm 29.9$ ∗∗
WiKG $24.1\pm 33.9$ ∗∗ $62.1\pm 26.4$ ∗∗ $\mathbf{82.8\pm 17.8}$ $0.68$
5.2.1 Stratified Analysis
To assess performance across lesion scales, we extracted connected lesion components and corresponding predicted heatmaps from the entire test set. Lesions were grouped into three size-based bins based on pixel area: Small ($10^{6}$ pixels), Moderate ($10^{6}$$10^{7}$ pixels), and Large ($10^{7}$ pixels). These thresholds were chosen based on the empirical area distribution in the dataset to reflect clinically meaningful differences. The bin counts are: $1415$ (Small), $135$ (Moderate), and $62$ (Large), showing that small lesions dominate the dataset and represent a particularly challenging and clinically important subgroup. Table 4 reports the Dice scores of INSIGHT and comparison models scores across different lesion sizes using both UNI and Virchow 2 backbones, along with one-tailed permutation test results (10,000 iterations) comparing INSIGHT to each baseline. INSIGHT consistently outperforms all baselines on Small and Moderate lesions ($p<0.01$), which are especially critical in clinical diagnosis due to their subtle presentation. For Large lesions, INSIGHT performs comparably or better than baselines. These findings highlight INSIGHT’s robustness across lesion scales and its particular advantage in detecting small, subtle pathological regions often missed by conventional MIL-based methods.
@@ 5.3 Qualitative Analysis
We visualize and compare heatmaps generated by INSIGHT and baseline methods in Fig. 4. Unlike baseline methods, which often rely on post-hoc adjustments, INSIGHT directly integrates interpretability into its architecture, mapping each patch’s probability to the heatmap. This eliminates the need for additional calibration steps, producing well-calibrated, fine-grained heatmaps that effectively highlight diagnostically significant regions. Additional visualizations are provided in Appendix B, which further illustrate the superior interpretability and precision of INSIGHT’s heatmaps across both WSIs and CT datasets.
Figure 4: Qualitative comparison of heatmaps generated by INSIGHT (ours) and baseline methods using UNI. For CAMELYON16, we compare lesion detection heatmaps generated by (a) our INSIGHT model and (b) the baseline methods. Additionally, we present (c) heatmaps for MosMed and (d) heatmaps for BRACS. Ground-truth areas are outlined in green. Note that BRACS does not provide detailed annotations, so ground truth is not displayed in panel (d). Additional visualizations can be found in the Appendix (Sec. B) for further reference.
@@
  • WiKG: Although WiKG captures interactions between patches through dynamic graph representation, its heatmaps occasionally exhibit overactivation in non-lesion areas. This impact boundary delineation and hurt interpretability in clinical applications.
Figure 5: Qualitative comparison of heatmaps generated by INSIGHT using different foundation models. We compare lesion detection heatmaps generated from (a) the original slide, (b) the ground truth in binary (c) INSIGHT with UNI, and (d) INSIGHT with Virchow 2. Yellow box region highlights differences in boundary sharpness and focus.
Impact of Foundation Model. To further evaluate the generalizability of INSIGHT across different foundation models, we compare heatmaps generated using Virchow 2 with those from the default UNI backbone. As shown in Fig. 5, both foundational models enable accurate localization of lesion regions; however, Virchow 2 produces slightly more refined boundaries in certain zoomed-in regions, suggesting that foundation model choice can affect visual precision. Despite this, INSIGHT’s overall interpretability and region alignment remain robust across backbones.
5.4 Comparison with Grad-CAM
To demonstrate the superiority of INSIGHT’s built-in heatmaps over post-hoc methods, we compare them with Grad-CAM on the MosMed and CAMELYON16 datasets (Table 5). Unlike Grad-CAM, which requires additional backward passes and is sensitive to noisy gradients, INSIGHT generates interpretable heatmaps directly during forward inference without extra computation. The results show that INSIGHT achieves significantly higher Dice scores across all datasets and foundation models, indicating more accurate and fine-grained localization of diagnostically relevant regions. This intrinsic explainability ensures better alignment with clinical expectations while avoiding the instability and coarse resolution often observed in Grad-CAM.